OpenStudy (anonymous):

Given the linear equation x – y = 5, find the slope of its graph. A. –5 B. –1 C. 1 D. 5

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@d3Xter

OpenStudy (anonymous):

x=y+5 subtract 5 from both sides to rearrange to: y=x-5 m = 1 (gradient) c = -5 (where the graph line intersects the y axis) (0,-5) is a point on the line to plot the graph more accurately u ideally need another 2 points. use the equation: y=x-5 to get another 2 points y=5-5 y=0 (when x = 5) (0,5) is a point on the line y=3-5 y= -2 (when x = 3) (3,-2) is a point on the line you have three points (0,-5), (5,0) and (3,-2). Draw an extended line through the middle of all three points and you have graphed x=y+5

OpenStudy (anonymous):

Your graph should look like this: http://www.wolframalpha.com/input/?i=graph+x%3Dy%2B5

OpenStudy (anonymous):

C.1

OpenStudy (anonymous):

^

OpenStudy (anonymous):

The answer was right there, and you cared less to see it. How do you learn that way?

OpenStudy (anonymous):

Uw anyways :)

OpenStudy (anonymous):

Very unethical @ifrimpanainte !